# An Ecosystem for Ontology Interoperability

### v5

@@ I Introduction
For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability between KGs. While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs. For example, it can be challenging to understand a KG snippet (Sydney, travelsTo, Sydney) alone, but with a class schema expression (Person, travelsTo, Place), one can infer that the subject “Sydney” refers to a person, while the object “Sydney” refers to the place being visited. The triple (Sydney, travelsTo, Sydney) describes an event in which the American actress Sydney Sweeney travels to the city of Sydney [1].
@@
Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain of application. There is no guarantee that ontologies in the same domain agree on the conceptualisation. Therefore, ontology interoperability has become a higher-level issue that affects information sharing between ontologies. For example, it is common for ontologies to encompass a range of different concepts, such as describing a person using their basic profiles (name, age, and gender), social relationships (parent-child and friendship), and professional roles (student and employee). Moreover, the terminology used in ontologies can be different. It is also common for one ontology to use “first name” and “last name”, but another ontology uses “given name” and “family name” to describe a person’s name.
@@
In this paper, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs). We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
II Related Work
II-A Ontology Design Patterns
The goal of ontology design patterns (ODPs) is to identify common ontology fragments across different ontologies [2]. Fig. 1 illustrates that ODPs use a star-based matching schema. Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies.
Figure 1: Ontology design patterns (ODPs).
@@ II-A Ontology Design Patterns
However, ODPs also have drawbacks. Firstly, ODPs are conceptual models and do not contain details of the actual class or property. This means that it is not possible to control the terminologies, and ontology creators are still free to use their own choices. Secondly, ODPs are commonly handcrafted on the basis of existing ontologies. They are static and tend not to capture changes over time. Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
II-B Ontology Matching and Versioning
Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies [3] or different versions of the same ontology [4]. Fig. 2 shows that OM&OV conform to a pairwise matching schema. This means that each mapping is a pairwise process to find the correspondences between two ontologies.
Figure 2: Ontology matching and versioning (OM&OV).
@@ II-B Ontology Matching and Versioning
OM&OV are typically resource-intensive. The complexity of pairwise mapping is the product of the number of entities in the two ontologies. The pairwise mapping process also needs to be repeated multiple times if multiple ontologies exist. Moreover, OM&OV tasks focus on either conceptual-level matching (i.e. TBox-to-TBox matching) or instance-level matching (i.e. ABox-to-ABox matching), and less attention has been paid to the compliance between conceptual models and data instances (i.e. TBox-to-ABox matching).
II-C Data-Driven Ontology Validation
Data-driven ontology validation (DOVA) uses KGs to validate the effectiveness of the ontology. Fig. 3 shows that DOVA complies with a star-based matching schema. Whenever ontologies comply with KG(s), their reference mappings can be derived from different ontology classes or properties pointing to the same KG entities.
Figure 3: Data-driven ontology validation (DOVA).
@@ II-C Data-Driven Ontology Validation
The introduction of ontology-compliant knowledge graphs (OCKGs) is a novel DOVA approach that uses a pay-as-you-go model to find the most cost-effective ontology fragment that matches the given KG(s) [5]. Building OCKGs is not easy. To determine which candidate ontology best fits the KG(s), the KG needs to be modelled for each candidate. This process essentially involves aligning the candidate ontologies, informed by the KG under study.
III An Ecosystem
Fig. 4 illustrates an overview of an ecosystem for ontology interoperability. Similar to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and data-driven ontology validation (DOVA) in different phases of the ontology engineering life cycle to achieve interoperability.
Figure 4: An overview of the ecosystem.
@@ III An Ecosystem
The design phase is responsible for understanding the requirements and creating a blueprint or prototype. In ontology engineering, these requirements are collected by competency questions. ODPs can serve as a skeleton to create conceptual models based on competency questions. The develop phase includes turning the prototype into a full product and applying a unit test. In ontology engineering, this phase is referred to as ontology construction and evaluation. OM&OV can be used to evaluate the level of interoperability between two ontologies, or between an ontology and its previous version. The deploy phase usually refers to implementing the product in the real world. In ontology engineering, we refer to this as applying the ontology to actual KGs. We can utilise OCKGs to comprehend the interoperability of ontologies within each KG or a group of KGs, thereby enhancing the design of the ontology in the next cycle.
IV A Case Study in the Building Domain
@@
IV-A Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is commonly recommended to reuse pre-existing ODPs. Fig. 5 illustrates that four classical ODPs can be used in this case to model the example of sensor observation from different perspectives. While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown below.
@@ IV-A Design Phase
Figure 5: Modelling using different ODPs.
On the other hand, differences in ontology characteristics, such as those listed below, need to be addressed in practice.
@@
IV-B Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and the original SSN [8, 12] (also called the old SSN in the W3C standard [13]). As illustrated in Fig. 6, SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to identify aligned concepts across two ontologies. SOSA and an updated SSN were co-developed, with a published alignment between them. Sosa was more of a module of the updated SSN (for which there were other modules too). If we consider SOSA to be an updated and reduced version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV can be unified with minor modifications [14]. OM&OV tasks can be classified into two main categories.
@@ IV-B Develop Phase
IV-C Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it is hard to determine which ontology to use without applying the real-world KGs. Some concepts used in a KG may not have the corresponding classes and properties in the ontology. Therefore, we apply these two ontologies in the KGs and construct the OCKGs for evaluation. Fig. 7 shows the OCKGs based on SOSA and SSN. We can see that both ontologies can capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the given KG.
@@ IV-C Deploy Phase
    We use sosa:hasFeatureOfInterest to replace ssn:FeatureOfInterest. Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we distinguish them to avoid confusion.
Fig. 8 shows the updated ontology and the OCKG built on the updated ontology. Apparently, the updated ontology is better suited to the provided KG instance. In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times. It is also possible to choose other alternative ODPs. For example, the class “BuildingSpace” in the Measurement Pattern is more suitable to describe a room, and the property “hasLocation/isPartOf” in the Event Pattern is more appropriate to describe the link between the sensor and the room.
@@ V Contributions
    OM&OV: We develop a software system for OM and OV. Agent-OM [17] uses state-of-the-art LLM agents that leverage its novel infrastructure to deliver improved OM performance in a fully-automated and domain-independent manner. Agent-OV [14] evolves from Agent-OM and explicitly targets OV tasks, including new task formulation and evaluation, with a cross-reference mechanism that compiles OM tasks. Agent-OM is used to create an LLM hallucination dataset [18] for the LLM leaderboard and LLM fine-tuning for OM. Agent-OM also participate in the Ontology Alignment Evaluation Initiative (OAEI) [19] 2025 campaign.
@@
    DOVA: We introduce OCKGs and define three subtasks under OCKGs: ontology compliance within KG, ontology compliance over KGs, and pattern-based compliance. Examples in the building domain showcase the efficiency of using OCKGs in ontology validation and decision-making [5].
VI Discussion & Conclusion
Ontology interoperability has been a long-standing issue studied in the Semantic Web community. Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings. Data-driven ontology validation aims to explore the interoperability between an ontology and its complementary KGs.
@@ VI Discussion & Conclusion
While there is no preference for one technique over another, these three techniques play different roles in an ecosystem for ontology interoperability. We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and using DOVA in the deploy phase could significantly improve interoperability throughout the ontology engineering life cycle.
Supplemental Material Statement
The source code, data, and/or other artifacts used in this study are available at https://github.com/qzc438-research/ontology-interoperability.
Acknowledgment

### v6

# ISWC 2026 Companion Volume, October 25-29, 2026, Bari, Italy

@@ 1 Problem Statement
For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability among KGs. While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs. For example, it can be challenging to understand a KG snippet (Sydney, travelsTo, Sydney) alone, but with a class schema expression (Person, travelsTo, Place), one can infer that the subject “Sydney” refers to a person, while the object “Sydney” refers to the place being visited. The triple (Sydney, travelsTo, Sydney) describes an event in which the American actress Sydney Sweeney travels to the city of Sydney [1].
@@
Ontologies are crafted by domain experts (in collaboration with the ontologist) based on their understanding of the domain of application. There is no guarantee that ontologies in the same domain share the same conceptualisation. Therefore, ontology interoperability has become a higher-level issue that affects the sharing and reuse of ontologies. Ontologies can encompass different entities to describe the same concepts. For example, describing a person can use their basic profiles (name, age, and gender), social relationships (parent-child and friendship), or professional roles (student and employee). The terminology used in the ontologies may also vary. One ontology can use “first name” and “last name” to describe a person’s full name, and another ontology can use “given name” and “family name” instead.
@@
To tackle this challenge, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and data-driven ontology validation (DOVA). We analyse their pros and cons and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to demonstrate how these three techniques can foster strengths and circumvent weaknesses throughout the ontology engineering life cycle.
2 Related Work
2.1 Ontology Design Patterns (ODPs)
@@
Ontology design patterns (ODPs) identify common ontology fragments across different ontologies [2]. ODPs use a star-based matching schema. Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies (see Figure 1). ODPs have some limitations. Firstly, ODPs are conceptual models and do not contain details of the actual class or property. This means that it is not possible to control terminology, and ontology creators remain free to make their own choices. Secondly, ODPs are commonly handcrafted based on existing ontologies. They are static and tend not to capture changes over time. Thirdly, the choice of ODPs is entirely reliant on one’s experience and preferences; the chosen ODP may be overwhelming, overlapping, or scarce for the implemented KG.
2.2 Ontology Matching and Versioning (OM&OV)
@@
Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies [3] or different versions of the same ontology [4]. OM&OV conform to a pairwise matching schema. This means that each mapping is a pairwise process to find the correspondences between two ontologies (see Figure 1). OM&OV are typically resource-intensive. The complexity of pairwise mapping is the product of the number of entities in the two ontologies. The pairwise mapping process also needs to be repeated multiple times if multiple ontologies exist. Moreover, OM&OV tasks focus on either conceptual-level matching (i.e. TBox-to-TBox matching) or instance-level matching (i.e. ABox-to-ABox matching), and less attention has been paid to the compliance between conceptual models and data instances (i.e. TBox-to-ABox matching).
2.3 Data-Driven Ontology Validation (DOVA)
@@
Data-driven ontology validation (DOVA) uses KGs to validate the effectiveness of the ontology. DOVA complies with a star-based matching schema. Whenever ontologies comply with KG(s), their reference mappings can be derived from different ontology classes or properties that point to the same KG entities (see Figure 1). The introduction of ontology-compliant knowledge graphs (OCKGs) is a novel DOVA approach that uses a pay-as-you-go model to find the most cost-effective ontology fragment that matches the given KG(s) [5]. Building OCKGs is not easy. To determine which candidate ontology best fits the KG(s), the KG needs to be modelled for each candidate ontology. This process essentially involves aligning the candidate ontologies, informed by the KG under study.
Figure 1: Illustration of different matching schema used in ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and data-driven ontology validation (DOVA).
3 An Ecosystem for Ontology Interoperability
@@
Figure 2 illustrates an overview of the ecosystem for ontology interoperability. Similar to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ODPs, OM&OV, and DOVA in different phases. (1) The design phase is responsible for understanding the requirements and creating a blueprint or prototype. In ontology engineering, these requirements are collected by competency questions. ODPs can serve as a skeleton to create conceptual models based on competency questions. (2) The develop phase includes turning the prototype into a full product and applying a unit test. In ontology engineering, this phase is referred to as ontology construction and evaluation. OM&OV can be used to evaluate the level of interoperability between two ontologies, or between an ontology and its previous version. (3) The deploy phase usually refers to implementing the product in the real world. In ontology engineering, we refer to this as applying the ontology to actual KGs. We can utilise OCKGs to assess the interoperability of ontologies within a single KG or across a set of KGs, thereby enhancing the design of the ontology in the next cycle.
Figure 2: An overview of the ecosystem. We propose to use ODPs in the design phase, OM&OV in the develop phase, and DOVA in the deploy phase to achieve better ontology interoperability.
4 A Case Study in the Building Domain
@@
4.1 Ontology Design Patterns (ODPs) in the Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is commonly recommended to reuse pre-existing ODPs. Figure 3 illustrates that four classical ODPs can be used in this case to model the example of sensor observation from different perspectives.
@@ 4.1 Ontology Design Patterns (ODPs) in the Design Phase
While all four of these ODPs can effectively capture the main concepts, significant differences in perspective need to be considered, as shown below.
@@
Figure 3: Modelling sensor observation using different ODPs.
4.2 Ontology Matching and Versioning (OM&OV) in the Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and the original SSN [8, 12] (also called the old SSN in the W3C standard [13]). As illustrated in Figure 4, SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to identify aligned concepts across two ontologies. SOSA and an updated SSN were co-developed, with a published alignment between them. Sosa was more of a module of the updated SSN (for which there were other modules too). If we consider SOSA to be an updated and reduced version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV can be unified with minor modifications [14]. OM&OV tasks can be classified into two main categories.
@@ 4.2 Ontology Matching and Versioning (OM&OV) in the Develop Phase
4.3 Data-Driven Ontology Validation (DOVA) in the Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it is hard to determine which ontology to use without applying the real-world KGs. Some concepts used in a KG may not have the corresponding classes and properties in the ontology. Therefore, we apply these two ontologies in the KGs and construct the OCKGs for evaluation. Figure 5 shows the OCKGs based on SOSA and SSN. We can see that both ontologies can capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the given KG.
@@ 4.3 Data-Driven Ontology Validation (DOVA) in the Deploy Phase
    We use sosa:hasFeatureOfInterest to replace ssn:FeatureOfInterest. Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we add the verb “has” to distinguish the property from the class.
Figure 6 shows the updated ontology and the OCKG built on the updated ontology. Apparently, the updated ontology is better suited to the provided KG instance. In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times. It is also possible to choose other alternative ODPs. For example, the class “BuildingSpace” in the Measurement Pattern is more suitable to describe a room, and the property “hasLocation/isPartOf” in the Event Pattern is more appropriate to describe the link between the sensor and the room.
@@ 5 Contributions Made So Far
    OM&OV: We study the interlink between modern LLMs and traditional OM systems [17]. We develop a novel software system for OM and OV. Agent-OM [18] uses state-of-the-art LLM agents that leverage its novel infrastructure to deliver improved OM performance in a fully-automated and domain-independent manner. Agent-OV [14] evolves from Agent-OM and explicitly targets OV tasks, including new task formulation and evaluation, with a cross-reference mechanism that compiles OM tasks. Agent-OM is used to create an LLM hallucination dataset [19] for the LLM leaderboard and LLM fine-tuning for OM. Agent-OM also participates in the Ontology Alignment Evaluation Initiative (OAEI) [20] 2025 campaign and showcases its generative capability in handling OM tasks at different levels of granularities [21].
@@
    DOVA: We introduce OCKGs and define three subtasks under OCKGs: ontology compliance within KG, ontology compliance over KGs, and pattern-based compliance. Examples in the building domain demonstrate the process of constraining OCKGs and showcase the efficiency of OCKGs in facilitating ontology validation and decision-making [5].
6 Conclusion and Future Work
@@
Ontology interoperability has been a long-standing issue studied in the Semantic Web community. Since the establishment of the International Semantic Web Conference (ISWC), workshops on ontology design and patterns (WOP) and ontology matching (OM) have been the two longest-running workshop proceedings. Data-driven ontology validation, which focuses on exploring the interoperability between an ontology and its complementary KGs, has received increasing attention in recent years. While there is no preference for one technique over another, these three techniques play different roles in an ecosystem for ontology interoperability. We believe that employing ODPs in the design phase, utilising OM&OV in the develop phase, and using DOVA in the deploy phase could significantly improve ontology interoperability throughout the ontology engineering life cycle.
Acknowledgements.